OpenStudy (anonymous):

Solve the system: y = 2x^2 + 2x - 1 y = -x^2 -x +5

OpenStudy (anonymous):

This is what I have so far: -x^2 - x + 5 = 2x^2 + 2x -1 0 = x^2 + x - 5 + 2x^2 + 2x- 1 0 = 3x^2+ 3x - 6

OpenStudy (anonymous):

do i just do the quadratic formula?

OpenStudy (aum):

Factor the 3 out first.

OpenStudy (aum):

3x^2+ 3x - 6 = 0 3(x^2 + x - 2) = 0 x^2 + x - 2 = 0 x^2 + 2x - x - 2 = 0 x(x + 2) - 1(x + 2) = 0 (x+2)(x-1) = 0 x = -2 or 1.

OpenStudy (anonymous):

wait where did the 2x come from?

OpenStudy (aum):

It is a way to factor a quadratic equation. The middle term x can be written as 2x - x

OpenStudy (anonymous):

ah thank you so much! :)

OpenStudy (aum):

You are welcome.